Anyone know how to solve this problem? No one has provided the right answer yet :/

Anyone know how to solve this problem? No one has
Answer #1

F_{x}= M\, a_{x}

a_{x}= \frac{10 N}{0.26 Kg}sin\left ( \frac{2\pi \, t}{4.0\, s} \right )

a_{x}= 38.46\, \frac{m}{s^{2}}sin\left ( \frac{2\pi \, t}{4.0\, s} \right )

\frac{dV_{x}}{dt} = 38.46\, \frac{m}{s^{2}}sin\left ( \frac{2\pi \, t}{4.0\, s} \right )

integrating the above equation we get,

V_{x}\left ( t \right ) = -\,\frac{4}{2\pi } 38.46 \frac{m}{s}cos\left ( \frac{2\pi \, t}{4.0\, s} \right )+\frac{4}{2\pi } 38.46\frac{m}{s}

V_{x}\left ( 2 \right ) = -\,\frac{4}{2\pi } 38.46 \frac{m}{s}cos\left ( \frac{2\pi \, 2}{4.0\, s} \right )+\frac{4}{2\pi } 38.46\frac{m}{s}

V_{x}\left ( 2 \right ) = 48.84 \frac{m}{s}